You are given a pendulum composed of a 0.050 kg mass on the end of a 0.40 m long massless string. If the pendulum is moved 30° from the vertical and given an initial speed of 0.29 m/s tangent to the support string and away from the vertical, how much higher relative to the release point of the pendulum swing will the pendulum rise?
Answer #1

Using Energy conservation:

KEi + PEi = KEf + PEf

KEi = (1/2)*m*Vi^2

m = mass of pendulum = 0.050 kg

Vi = Initial speed by which pendulum is launched = 0.29 m/sec

PEi = 0, assuming the point from which pendulum is released as a reference point

KEf = 0, since at max height, till which pendulum will rise, speed will be zero

PEf = m*g*h

h = height of pendulum from reference point

Using given values:

(1/2)*m*Vi^2 + 0 = 0 + m*g*h

h = Vi^2/(2*g)

h = 0.29^2/(2*9.81)

h = 4.29*10^-3 m

So relative to the release point pendulum will rise 4.29*10^-3 m

Please Upvote.